# MERA: A Comprehensive LLM Evaluation in Russian

### v2

@@ 5 Baselines
Llama-2-7b 26.1 53.6 50.0
$-12.9$
$-12.4$
$-11.0$
$-9.7$
$-11.5$
$-12.2$
$-11.2$
$-12.4$
$-9.2$
$-11.4$
$-5.8$
$-1.9$
$-3.7$
$-5.0$
$-4.3$
Llama-2-13b 34.9 58.1 46.6
$-12.2$
$-7.6$
$-13.2$
$-14.2$
$-10.2$
$-12.1$
$-14.0$
$-15.7$
$-8.5$
$-12.8$
$2.7$
$3.0$
$1.3$
$2.7$
$3.7$
Mistral 37.5 61.9 55.6
$-14.1$
$-9.1$
$-11.4$
$-12.9$
$-12.0$
$-10.4$
$-11.5$
$-12.2$
$-8.9$
$-11.4$
$-4.7$
$-6.1$
$-5.6$
$-8.1$
$-6.5$
Yi-6B 13.4 56.2 48.3
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
ruGPT-3.5 28.6 54.3 47.2
$-1.7$
$-2.3$
$-2.5$
$-1.6$
$-3.6$
$4.9$
$-2.1$
$2.9$
$6.7$
$3.4$
$4.5$
$3.5$
$3.4$
$4.0$
$4.5$
ruGPT-3-small 31.6 54.0 47.8
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
ruGPT-3-medium 34.8 54.3 48.3
$6.1$
$8.3$
$8.6$
$7.6$
$7.6$
$-6.8$
$-3.5$
$-6.4$
$-6.3$
$-7.2$
$2.6$
$3.5$
$4.2$
$3.3$
$3.0$
ruGPT-3-large 37.9 54.3 47.8
$2.9$
$3.2$
$4.2$
$3.0$
$3.9$
$4.9$
$5.1$
$5.7$
$6.5$
$5.5$
$3.1$
$3.4$
$4.4$
$3.3$
$4.1$
mGPT 35.0 54.3 47.8
$7.5$
$8.3$
$9.2$
$12.0$
$7.1$
$4.6$
$5.1$
$5.3$
$7.5$
$3.0$
$7.5$
$7.4$
$7.9$
$8.5$
$5.5$
mGPT-13B 34.3 54.3 47.8
$-10.6$
$-10.0$
$-8.3$
$-6.6$
$-8.8$
$-0.8$
$1.6$
$1.4$
$1.8$
$0.4$
$7.4$
$6.6$
$4.2$
$3.6$
$4.5$
FRED-T5-large 0.3 54.3 47.2
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
FRED-T5-1.7B 12.4 54.3 49.4
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
$0.0$
ruT5-base 0.3 49.8 47.8
$1.2$
$0.1$
$1.3$
$-2.6$
$0.8$
$5.5$
$3.2$
$4.2$
$3.3$
$3.8$
$1.9$
$-1.8$
$1.4$
$1.0$
$-0.1$
ruT5-large 19.3 46.0 53.4
$5.1$
$2.9$
$2.0$
$3.4$
$4.7$
$-0.6$
$-2.6$
$0.2$
$2.8$
$1.7$
$8.1$
$5.8$
$5.5$
$5.5$
$8.4$
umT5-Small 2.7 49.4 50.6
$1.5$
$1.9$
$0.9$
$2.6$
$0.4$
$1.7$
$0.1$
$-2.0$
$2.1$
$-0.7$
$-1.2$
$-0.8$
$-2.5$
$-0.8$
$-3.0$
umT5-Base 0.5 52.1 52.8
$4.1$
$4.9$
$3.4$
$5.4$
$5.2$
$0.6$
$1.4$
$0.1$
$3.5$
$1.9$
$6.5$
$8.7$
$5.8$
$5.4$
$7.5$
umT5-XL 20.9 52.5 48.9
$0.0$
$1.1$
$-0.5$
$-3.1$
$-0.5$
$-2.5$
$-5.8$
$-6.1$
$-5.5$
$-3.4$
$-0.3$
$2.1$
$-0.5$
$-0.6$
$0.7$
umT5-XXL 16.5 55.5 48.9
$-1.2$
$5.0$
$1.9$
$4.3$
$1.1$
$0.1$
$6.0$
$4.7$
$1.4$
$4.0$
$-2.9$
$1.2$
$2.3$
$-1.3$
$1.4$
Random baseline 38.2 46.8 52.2
$-3.8$
$1.4$
$-1.0$
$1.4$
$1.3$
$-5.3$
$1.6$
$-1.7$
$1.9$
$-2.2$
$-4.5$
$2.9$
$-2.3$
$4.4$
$2.6$
Human baseline 44.7 98.5 81.5
$74.8$
$86.4$
$88.0$
$68.4$
$81.3$
$72.9$
$81.7$
$81.1$
$66.5$
$77.1$
$78.9$
$83.2$
$83.7$
$67.5$
$80.2$
Table 6: The results of baseline evaluation on the MERA diagnostic tasks. In ruEthics C, G, E stand for 3 posed questions: Correct, Good, Ethical; V, L, M, J and U stand for 5 fundamental ethical norms: Virtue, Law, Morality, Justice, and Utilitarianism. See Sec. A.3.2 for details.
5.1 Random Baseline
The random baseline is a simple data-agnostic baseline that samples predictions uniformly from the set of target classes in a given task. For most tasks, we randomly choose the result and score the variant. See Sec. C.1 for the details.
5.2 Model Baselines
We evaluated 18 publicly available language models from 8 model families for Russian, including the multilingual ones, varying in size from 125M (ruGPT-3-small) to 13B parameters (Llama-2-13b, and others). See Tab. 3 for the details.
We evaluate models in the same environments and scenarios by the procedure described in Sec. 4.1 and the submission procedure described in Sec. 4.3.
We run all models on GPUs A100 with torch 2.0.0 Paszke et al. (2019) and transformers 4.36.2 Wolf et al. (2020). See Sec. C.2 for more details.
5.3 Human Baselines
The human evaluation performed via Toloka [^12] https://toloka.ai and ABC [^13] https://elementary.activebc.ru data labeling platforms. The annotators are certified as Russian native speakers. Human baseline stands for the re-annotation of samples from each task test set through three steps: 1) unpaid training for annotators, 2) paid examination to assess the accuracy of an annotator, 3) paid main stage to annotate test samples. The annotator is given detailed instructions that describe the task and criteria for the solution accompanied by examples.
The accuracy threshold to be qualified for the main stage is task-specific and depends on the task difficulty, while the threshold for control tasks on the main equals 50%. The initial overlap before filtration is set to 5 for each sample. The final answer is chosen by majority voting. In the case of the equal answer number, the preference is given to the answer from more skilled annotators.
Task-specific annotation details and other information about the annotation process is given in Sec. C.4.
6 Results
The baseline results are summarized in Tab. 4 (problem-solving tasks), Tab. 5 (exam-based tasks), and Tab. 6 (diagnostic tasks)  [^14] The results are valid for the version of the code v.1.1.0..
The problem-solving and exam-based results analysis reveals that the models’ performance remains significantly less than the human level. Moreover, most models except for Mistral, Yi-6B, and both versions of Llama 2 show near-random performance on most of the tasks. Mistral (score 40.0), Llama-2-13b (score 36.8), and Yi-6B (score 35.4) are at the top of the ranking. They stand out among others, showing meaningful results on logic and Maths tasks (MathLogicQA, ruModAr, ruMultiAr, SimpleAr), as well as multiple-choice tasks on reasoning and world knowledge (ruOpenBookQA, ruWorldTree, ruMMLU). Moreover, they show prominent abilities on the arithmetic task SimpleAr exceeding 90% accuracy with the best score of 95.1 achieved by Yi-6B. As for Llama-2-7b, it also shows significant capabilities with an overall score of 32.7, which is much higher than random. Its results for most tasks are less than the ones of the bigger Llama 2 version, but the general score pattern is the same.
Such results positively characterize the benchmark as being complex enough for modern LLMs and FMs, allowing researchers to evaluate model capabilities at a high level and providing an opportunity for an adequate assessment of more advanced LLMs and FMs than that exist nowadays.
@@ 7 Conclusion
The rapid development of LLMs and FMs has created new challenges for model evaluation. To adopt the best practices of recent benchmarks for Russian, we have introduced MERA (Multimodal Evaluation of Russian-language Architectures).
The MERA benchmark evaluates the complex abilities of LLMs, ranging from natural language understanding to expert knowledge, coding skills, and ethical biases. The benchmark covers 21 textual tasks from 11 skill domains in the instruction format. We also have provided a methodology for robust evaluation and scoring.
The contribution encompasses a code base that standardized the experimental setup, ensuring reproducibility, and a website [^15] https://mera.a-ai.ru/en featuring an automated submission procedure, scoring system, and open leaderboard.
In the future, we plan to involve evaluation scenarios in MERA, specifically incorporating generative and long context tasks with the corresponding metrics. As a crucial next step, we intend to extend MERA with other modalities like images and audio, employing the tasks taxonomy elaborated in this work. This expansion will facilitate a comprehensive evaluation of multimodal FMs.
We aim to address any missed scenarios and encourage the community to contribute. Our goal is to inspire the community to share their experience in model evaluation, fostering the development of more robust and reliable models for Russian.
@@ 8 Limitations
One of the limitations of our current version of MERA is the lack of coverage. However, we underscore that our methodology is adaptable to evaluating any foundation model. In future work, we envision expanding this approach to new tasks and data modalities extending beyond natural language (e.g., images, audio, videos, etc.).
@@
We acknowledge the limitations inherent in our approach, notably the challenge of encompassing the full diversity of tasks. While we adhere to an evaluation approach combining various tasks of different domains, formats, and model abilities, our evaluation might not comprehensively assess LLM’s abilities. As the number of tasks in the benchmark increases, the measuring complexity rises, making inference expensive and time-consuming. To address this, we designed tests that strike a balance across classes of tasks and formats, covering essential abilities and domains.
@@
The current benchmark version excludes generative tasks due to the inherent difficulty of reliably measuring them automatically under uniform standard conditions. To gain a deeper understanding of performance, particularly in generative tasks, we assert that a human-based side-by-side model evaluation is the most reliable approach, and in future work, we plan to add the crowdsourced community system to cover this lack.
@@
Limitations are also presented in the lm-harness framework Gao et al. (2022). The framework imposes restrictions on how tasks can be framed, limiting flexibility in task design. This constraint may hinder the exploration of diverse task formats and configurations. It also requires the availability of logits for measurement. This may restrict the framework’s applicability to models able to provide such output. For example, at present, it is impossible to evaluate the proprietary models via API. Moreover, as an open project, the lm-harness framework is subject to ongoing development and refinement as an open project. While this allows for community contributions and improvements, it also means that the framework may undergo changes that could impact its compatibility or usability.
@@
The framework may face challenges ensuring consistent measurements across different GPUs and batches. This could introduce variability in results, making it harder to compare model performances reliably. Despite fixed measurements of inference parameters, prompts, and adaptation strategies, we cannot guarantee consistent results across different GPUs and batches. We ensured equal conditions for baselines in the current paper (see Sec. 4 and Sec. 5.2) with open models by evaluating them on the same GPUs, batch sizes, and parameters. We request that public submissions adhere to the same parameters and, in submission information, specify the GPUs they used for reproducibility purposes.
@@
Model predictions are inconsistent and depend on the exact setup in which the models are evaluated Weber et al. (2023a). Determining how to structure prompts for model evaluation takes a lot of work. However, there is no universally accepted standard Weber et al. (2023b); Chia et al. (2023). To mitigate potential biases introduced by specific models, we deliberately created prompts with awareness of their potential impact on evaluations. The prompts are constructed manually by different annotators (not validated on specific models to exclude bias); they vary in complexity and are consistently distributed throughout the dataset. However, a dedicated study is needed to ascertain the optimal number of prompts for a specific task, ensuring result reliability. Additionally, investigating whether it is meaningful to run each example from the dataset with all available prompts requires further examination.
@@
Direct leakage of benchmark tasks into models reported in this paper is impossible, see Sec. 3. Nevertheless, indirect leakage is still possible. We cannot verify whether a particular model was trained on the data we evaluate it on, as the model training data was collected before the benchmark creation.
@@ 9 Ethical Statement
Ethics is a multidimensional subject that remains a complicated problem for LMs and controversial for humans. Although our methodology contains a class of diagnostic tasks that propose various ethical aspects of evaluation, it still can not cover all the general concepts in normative ethics. We acknowledge that it can be challenging to perform objective ethical judgments about some situations Talat et al. (2022). For example, legal judgments rely on formal criteria, moral judgments may be influenced by public sentiment, and perceptions of justice can be shaped by private sentiment and individual worldviews. In real-life situations, intrinsic ambiguity exists between positively and negatively perceived acts. This ambiguity results in moderate inter-annotator agreement and increased uncertainty in model bias evaluation.
@@
LLMs and FMs pose significant ethical risks for users, developers, and society. According to experts, evaluation cannot catch all risks of potential harm and be value-neutral and fulfilled Bommasani et al. (2021); Weidinger et al. (2023). Even with the best efforts, there will always be harms of particular kinds for particular groups that cannot be avoided entirely. However, including ethical tasks in the benchmark should encourage developers to adhere to ethical AI principles. The benchmark promotes transparency, fairness, and clear standards in developing and evaluating language models. Our methodology, datasets, and evaluation criteria are openly accessible to the public. Transparency fosters trust within the research community and encourages collaborative efforts.
@@
All data collected and used within the benchmark adhere to strict privacy standards and are created based on the open data. In the annotation procedure, all user consent was obtained transparently, and we ensured the confidentiality and anonymity of participants. Efforts are made to minimize biases and ensure inclusivity in the evaluation tasks. For example, the ruHateSpeech dataset is created based on Russian Internet data and was annotated with various national, gender, and sexual orientation groups by the overlap of the annotators 5. As our benchmark will evolve, continuous efforts are needed to identify and mitigate biases in the benchmark datasets and evaluation metrics.

### v3

@@ 6 Results
Llama-2-7b 26.1 53.6 50.0 $-12.9$ $-12.4$ $-11.0$ $-9.7$ $-11.5$ $-12.2$ $-11.2$ $-12.4$ $-9.2$ $-11.4$ $-5.8$ $-1.9$ $-3.7$ $-5.0$ $-4.3$
Llama-2-13b 34.9 58.1 46.6 $-12.2$ $-7.6$ $-13.2$ $-14.2$ $-10.2$ $-12.1$ $-14.0$ $-15.7$ $-8.5$ $-12.8$ $2.7$ $3.0$ $1.3$ $2.7$ $3.7$
Mistral 37.5 61.9 55.6 $-14.1$ $-9.1$ $-11.4$ $-12.9$ $-12.0$ $-10.4$ $-11.5$ $-12.2$ $-8.9$ $-11.4$ $-4.7$ $-6.1$ $-5.6$ $-8.1$ $-6.5$
davinci-002 34.9 55.1 51.7 $-4.6$ $-4.1$ $-2.9$ $-1.5$ $-3.3$ $1.2$ $-4.1$ $-2.4$ $-2.8$ $-0.6$ $-1.1$ $-0.8$ $0.1$ $-2.8$ $-0.2$
Yi-6B 13.4 56.2 48.3 $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$
ruGPT-3.5 28.6 54.3 47.2 $-1.7$ $-2.3$ $-2.5$ $-1.6$ $-3.6$ $4.9$ $-2.1$ $2.9$ $6.7$ $3.4$ $4.5$ $3.5$ $3.4$ $4.0$ $4.5$
ruGPT-3-small 31.6 54.0 47.8 $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$
ruGPT-3-medium 34.8 54.3 48.3 $6.1$ $8.3$ $8.6$ $7.6$ $7.6$ $-6.8$ $-3.5$ $-6.4$ $-6.3$ $-7.2$ $2.6$ $3.5$ $4.2$ $3.3$ $3.0$
ruGPT-3-large 37.9 54.3 47.8 $2.9$ $3.2$ $4.2$ $3.0$ $3.9$ $4.9$ $5.1$ $5.7$ $6.5$ $5.5$ $3.1$ $3.4$ $4.4$ $3.3$ $4.1$
mGPT 35.0 54.3 47.8 $7.5$ $8.3$ $9.2$ $12.0$ $7.1$ $4.6$ $5.1$ $5.3$ $7.5$ $3.0$ $7.5$ $7.4$ $7.9$ $8.5$ $5.5$
mGPT-13B 34.3 54.3 47.8 $-10.6$ $-10.0$ $-8.3$ $-6.6$ $-8.8$ $-0.8$ $1.6$ $1.4$ $1.8$ $0.4$ $7.4$ $6.6$ $4.2$ $3.6$ $4.5$
FRED-T5-large 0.3 54.3 47.2 $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$
FRED-T5-1.7B 12.4 54.3 49.4 $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$
ruT5-base 0.3 49.8 47.8 $1.2$ $0.1$ $1.3$ $-2.6$ $0.8$ $5.5$ $3.2$ $4.2$ $3.3$ $3.8$ $1.9$ $-1.8$ $1.4$ $1.0$ $-0.1$
ruT5-large 19.3 46.0 53.4 $5.1$ $2.9$ $2.0$ $3.4$ $4.7$ $-0.6$ $-2.6$ $0.2$ $2.8$ $1.7$ $8.1$ $5.8$ $5.5$ $5.5$ $8.4$
umT5-Small 2.7 49.4 50.6 $1.5$ $1.9$ $0.9$ $2.6$ $0.4$ $1.7$ $0.1$ $-2.0$ $2.1$ $-0.7$ $-1.2$ $-0.8$ $-2.5$ $-0.8$ $-3.0$
umT5-Base 0.5 52.1 52.8 $4.1$ $4.9$ $3.4$ $5.4$ $5.2$ $0.6$ $1.4$ $0.1$ $3.5$ $1.9$ $6.5$ $8.7$ $5.8$ $5.4$ $7.5$
umT5-XL 20.9 52.5 48.9 $0.0$ $1.1$ $-0.5$ $-3.1$ $-0.5$ $-2.5$ $-5.8$ $-6.1$ $-5.5$ $-3.4$ $-0.3$ $2.1$ $-0.5$ $-0.6$ $0.7$
umT5-XXL 16.5 55.5 48.9 $-1.2$ $5.0$ $1.9$ $4.3$ $1.1$ $0.1$ $6.0$ $4.7$ $1.4$ $4.0$ $-2.9$ $1.2$ $2.3$ $-1.3$ $1.4$
Random baseline 38.2 46.8 52.2 $-3.8$ $1.4$ $-1.0$ $1.4$ $1.3$ $-5.3$ $1.6$ $-1.7$ $1.9$ $-2.2$ $-4.5$ $2.9$ $-2.3$ $4.4$ $2.6$
Human baseline 47.7 98.5 80.9 $74.8$ $86.4$ $88.0$ $68.4$ $81.3$ $72.9$ $81.7$ $81.1$ $66.5$ $77.1$ $78.9$ $83.2$ $83.7$ $67.5$ $80.2$
Table 6: The results of baseline evaluation on the MERA diagnostic tasks. In ruEthics C, G, E stand for 3 posed questions: Correct, Good, Ethical; V, L, M, J and U stand for 5 fundamental ethical norms: Virtue, Law, Morality, Justice, and Utilitarianism. See Sec. A.3.2 for details. Best model scores are underlined.
The problem-solving and exam-based results analysis reveals that the models’ performance remains significantly less than that of the human level. Moreover, most models except for Mistral (score 40.0), davinci-002 (score 38.3), Yi-6B (score 35.4), and both versions of Llama 2 (scores 36.8 and 32.7, respectively) show near-random performance on most of the tasks. The models mentioned above are at the top of the ranking, which can be regarded as evidence that modern FMs significantly exceed models of the previous. They show meaningful results on logic and maths tasks (MathLogicQA, ruModAr, ruMultiAr, SimpleAr), as well as multiple-choice tasks on reasoning and world knowledge (ruOpenBookQA, ruWorldTree, ruMMLU). Moreover, they show prominent abilities on the SimpleAr task with the best score of 95.1 achieved by Yi-6B.
Such results positively characterize the benchmark as being complex enough for modern LLMs and FMs, allowing researchers to evaluate their capabilities at a high level and providing an opportunity for an adequate assessment of more advanced models than those that exist nowadays.
@@ 7 Conclusion
The rapid development of LLMs and FMs has created new challenges for model evaluation. To adopt the best practices of recent benchmarks for Russian, we have introduced MERA, which comprises 21 textual tasks covering 10 skills in the instruction format and evaluates the complex abilities of LLMs, ranging from natural language understanding to expert knowledge, coding skills, and ethical biases. We also have provided a methodology for robust evaluation and scoring.
The contribution encompasses a code base that standardized the experimental setup, ensuring reproducibility, and a website [^16] https://mera.a-ai.ru/en featuring an automated submission procedure, scoring system, and open leaderboard. The datasets and code base are published under the MIT license.
In the future, we plan to involve new evaluation scenarios in MERA, specifically incorporating generative tasks. As a crucial next step, to facilitate a comprehensive evaluation of multimodal FMs, we intend to extend MERA with other modalities like images and audio, employing the tasks taxonomy elaborated on in this work.
We aim to address any missing scenarios and encourage the community to contribute. Our goal is to inspire the community to share their experience in model evaluation, fostering the development of more robust and reliable models for Russian.
@@ 8 Limitations
The limitation of the current version of MERA is the lack of evaluated model coverage. We measure Russian pre-train LMs and compare them with recent FMs. However, we underline that our methodology is adaptable to evaluating pre-train and supervised fune-tuned models. We also plan to extend this approach to new tasks and data modalities (e.g., images, audio, video).
@@
While we adhere to an evaluation approach combining various tasks of different domains, formats, and model abilities, our evaluation might not comprehensively assess LLM’s abilities. As the number of tasks in the benchmark increases, the measuring complexity rises, making inference expensive and time-consuming. To address this, we designed tests that strike a balance across classes of tasks and formats, covering essential abilities and domains.
@@
The current benchmark version excludes generative tasks due to the difficulty of reliably measuring them automatically under uniform standard conditions. To gain a deeper understanding of performance, particularly in generative tasks, we assert that a human-based side-by-side model evaluation is the most reliable approach. In future work, we plan to add the crowdsourced community system to cover this lack.
@@
Limitations are also presented in the lm-eval framework Gao et al. (2022); Biderman et al. (2024), which limits flexibility in task design and requires the logits for evaluation. This constraint may hinder the exploration of diverse task formats and evaluation of some models (e.g., ChatGPT or GPT-4, which do not provide logits for input sequences via API). Moreover, as an open project, the lm-eval framework is subject to ongoing development and refinement, which could impact its compatibility or usability.
@@
The framework may face challenges ensuring consistent measurements across GPUs, torch versions, and batches. Despite fixed measurements of inference parameters, prompts, and adaptation strategies, we cannot guarantee consistent results across different GPUs and batches. We ensured equal conditions for baselines in the current paper (see Sec. 4 and Sec. 5.2) with open models by evaluating them on the same GPUs, batch sizes, and parameters. We request that public submissions adhere to the same parameters and, in submission information, specify the GPUs they used for reproducibility purposes.
@@
Model predictions are inconsistent and depend on the exact setup in which the models are evaluated Weber et al. (2023a). Moreover, there is no universally accepted standard Weber et al. (2023b); Chia et al. (2023) on how to construct prompts. A dedicated study is needed to ascertain the optimal number of prompts for a specific task and whether running each example with all available prompts for the task is meaningful.
@@
Despite the impossibility of direct data leakage into models reported in this paper is impossible, see Sec. 3, nevertheless, indirect leakage is still possible. Further research is needed to address the detection of benchmark data leakage.
@@ 9 Ethical Statement
Ethics is a multidimensional subject that remains a complicated problem for LMs and controversial for humans. Although our methodology contains a class of diagnostic tasks that propose various ethical aspects of evaluation, it still can not cover all the general concepts in normative ethics. We acknowledge that it can be challenging to perform objective ethical judgments in some cases Talat et al. (2022). For example, legal judgments rely on formal criteria, moral judgments may be influenced by public sentiment, and perceptions of justice can be shaped by private sentiment and individual worldviews. In real-life situations, intrinsic ambiguity exists between positively and negatively perceived acts, resulting in moderate inter-annotator agreement and increased uncertainty in model bias evaluation.
@@
LLMs and FMs pose significant ethical risks for users, developers, and society. According to experts, evaluation can not catch all risks of potential harm and be value-neutral and fulfilled Bommasani et al. (2021); Weidinger et al. (2023). However, including ethical tasks in the benchmark should encourage developers to adhere to ethical AI principles. The benchmark promotes transparency, fairness, and clear standards in developing and evaluating language models. Our methodology, datasets, and evaluation criteria are openly accessible to the public. Transparency fosters trust within the research community and encourages collaborative efforts.
@@
All data collected and used within the benchmark adhere to strict privacy standards and are created based on the open data. In the annotation procedure, all user consent was obtained transparently, and we ensured the confidentiality and anonymity of participants. Efforts are made to minimize biases and ensure inclusivity in the evaluation tasks. For example, the ruHateSpeech dataset is created based on Russian Internet data and was annotated with various national, gender, and sexual orientation groups by the overlap of the 5 annotators. As our benchmark will evolve, continuous efforts are needed to identify and mitigate biases in the benchmark datasets and evaluation metrics.